A store specializing in mountain bikes is to open in one of two malls. If the first mall is​ selected, the store anticipates a yearly profit of $ 675,000 if successful and a yearly loss of $ 225,000 otherwise. The probability of success is 1/2. If the second mall is​ selected, it is estimated that the yearly profit will be $ 450,000 if​ successful; otherwise, the annual loss will be $ 135,000. The probability of success at the second mall is 3/4.
a. What is the expected profit for the first​ mall? 
b. What is the expected profit for the second​ mall? 

Answer by Boreal(15235)

This would be half of the difference of $225000
E(X)=0.5*(675)-0.5*225=337.5-112.5 or $225 thousands or $225,000.
The second is 450*(3/4)-135(1/4)=$303.75 thousands or $303,750.